Number 8
Find the surface area of the bookends shown.
the first one has the numbers shown as 16.5cm 4cm 11cm 16cm. The second one shows 16cm 4 cm 11cm 16.5cm. With 5 sides total I beleive. I beleive that it is the bookends that are shown.
1
Expert's answer
2012-06-07T07:45:17-0400
1) Surface area is 2 areas of triangles with sides (16, 16.5, 4) and 3
rectangles built on these sides of height 11.
2) S(triangle) =
sqrt(p*(p-a)*(p-b)*(p-c))=sqrt(18.25*2.25*14.25*1.75)=32, where
p=(a+b+c)/2
3) S(sides)=(a+b+c)*h=(36.5)*11=401.5;
4) Surface
area=401.5+32+32=465.5